4. The company Crunchy Hippy Granola Inc. anticipates that the num- ber of new customers (Y) resulting from an intensive marketing cam- paign in the Chicago area will follow (approximately) a log-normal distribution with mean ?? 800 and variance ?2-512. It is known that X-lny follows a normal distribution where ? and ?2 denote the mean and variance of (a) Why should the lognormal distribution only be considered an ap- proximation here? (Consider the variable being measured and comment. (b) Find the mean and varia (c) Calculate P(Y ? 600), the probability that Hipster Granola, Inc (d) Find the 40th and 95th percentiles of Y nce of the normal random variable (Hint: it may be convenient here to work with CVY-?y/? will have at least 600 new customers in the Chicago area .
Answer #1

The log.normal distribution is the distribution that can flexibly takes any non-negative values. Alse, log-normal distributio

Now, find the mean of X as follows: σ2 expl μ+- 800 Squaring both sides,

expl μ+- (800) exp | 21 μ+- (800) exp(2μ + σ*) ny -ap(o)-1 exp (800) 12 exp (σ)-3824697 ơ-in (382.4697) σ = 5.95 Now, compute

5.95 800 = expl μ +-- In (800) = μ + 2.97 μ=668-297 = 3.71 Conclusion: The mean and variance of normal random variable Xis 3.

Compute the probability that Hipster Granola Inc. will have at least 600 new customers in Chicago area as follows: P(Y>600) 1

Conclusion: There are approximately 14% chances that Hipster Granola Inc. will have at least 600 new customers in Chicago are

-0.25 In(y)-3.71_-025 V5.95 5.95 In (y)-0.25x V5.95) +3.71 exp (3.1002) у 22.20 У~22 Conclusion: The 40h percentile of new cu

Using standard normal table, find the value of z with the probability nearest to 0.9, that is, 1.645. Therefore, P(z <1.645)

Conclusion: The 956 percentile of new customers is approximately 2259